OpenStudy (anonymous):

Simplify: (a2 + 2a − 8 ) − (5a − 5) (A) a2 − 7a − 13 (B) a2 − 3a − 13 (C) a2 − 7a − 3 (D) a2 − 3a − 3

OpenStudy (anonymous):

-*-=+

OpenStudy (anonymous):

(a2+2a-8-5a+5) =a2+2a-5a+5-8 =a2-3a-3